OpenStudy (anonymous):

Suppose that an object is traveling along a straight line. Suppose that its position function is given by f(t) = 5t^3 +11t where t is in seconds. Find the object's instantaneous acceleration at t=9 seconds. 182 210 244 270 305

OpenStudy (anonymous):

a is the second derivative of position with respect to time \[\frac {df} {dt}= 15t^{2}+11\]\[\frac {d^{2}f}{dt{^2}}=30t=30(9)=270\]